# Streaming algorithms for products of probabilities

### v1

@@ Abstract
We consider streaming algorithms for approximating a product of input probabilities up to multiplicative error of $1-\epsilon$. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(\log n+\log b-\log\epsilon)-\mathcal{O}(1)$, where $n$ is length of the input stream and $b$ is the bit length of the input numbers. This matches an upper bound from Alur et al. up to a constant multiplicative factor. Moreover, we consider the threshold problem, where it is asked whether the product of the input probabilities is below a given threshold. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(n\cdot b)$.
keywords:
streaming algorithms
category:
@@ 1 Introduction
Our first result is a lower bound that matches the upper bound $22\log n+\log b-\log\epsilon+$ 2 up to a constant factor. This is also true for randomized streaming algorithms that produce a $(1-\epsilon)$-approximation of the product $\prod_{1\leq i\leq n}q_{i}$ with high probability. Our lower bound only holds if $\log(1/\epsilon)\leq b<\log(1/(1-\epsilon))\cdot n/2$ holds. This is a reasonable setting. If we consider the approximation parameter $\epsilon$ as a small constant then the requirement becomes $\Omega(1)\leq b\leq\mathcal{O}(n)$. Since the length $n$ of the stream is usually very large, it is reasonable to assume $b\leq\mathcal{O}(n)$. In a realistic scenario one would even require $b\leq\mathcal{O}(\log n)$.
Our second result is a new lower bound for the threshold problem: the input consists again of a stream of probabilities $q_{1},q_{2},\ldots,q_{n}$, where $q_{i}=r_{i}/s_{i}$ and the integers $r_{i}$ and $s_{i}$ need at most $b$ bits. In addition, there is a threshold probability $t=r/s$, where $r$ and $s$ also need at most $b$ bits. The question is whether $\prod_{1\leq i\leq n}q_{i}<t$ holds. It was shown in [1, Theorem A.1] that every deterministic streaming algorithm for the corresponding problem, where $b=\log n$ and the $q_{i}$ and $t$ are not restricted to the interval $[0,1]$, has to store $\Omega(n)$ many bits. It was asked whether this lower bound also holds when all $q_{i}$ and $t$ are probabilities. We prove that this is indeed the case and moreover improve the lower bound from $\Omega(n)$ to $\Omega(n\cdot b)$ if $b\geq\Omega(\log n)$ for randomized streaming algorithms. Note that the threshold problem can be always solved in space $\Omega(n\cdot b)$ by naively computing the product $\prod_{1\leq i\leq n}q_{i}$. On the other hand, if $n$ is very large compared to $b$, then there is a better (deterministic) algorithm that only needs $\mathcal{O}(2^{b})$ bits; see Theorem 4.1.
In the final section we consider the product approximation problem for rational probabilities in the sliding window setting. Here, the goal is to maintain a $(1-\epsilon)$-approximate of the product of the last $m$ probabilities from the input stream. The parameter $m$ is called the window size. We provide a space lower bound of $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$ for randomized sliding window algorithms, where as before $b$ is the bit length of the input probabilities. Currently, we are not able to match this lower bound by an upper bound. In Section 5 we provide two incomparable upper bounds: $2mb$ (which is trivial) and $m\cdot(\log m+\log b-\log\epsilon)$.
@@ 2 Preliminaries
For a possibly infinite set $A$ we denote with $A^{*}$ the set of all finite words (or sequences) over the set $A$. We write $A^{\leq n}$ for the subset of words of length at most $n$. With $\log$ we always denote the logarithm to the base two. For a positive rational number $q=r/s$ with $r,s\in\mathbb{N}$ and $\mathsf{gcd}(r,s)=1$ we define the bit size of $q$ as $\|q\|=\max\{\lceil\log r\rceil,\lceil\log s\rceil\}$. In addition, we define $\|0\|=0$. Let =$\mathbb{P}=\{q\in\mathbb{Q}:0\leq q\leq 1\}$ be set of rational probabilities and $\mathbb{P}_{b}=\{q\in\mathbb{P}:\|q\|\leq b\}$ be the set of rational probabilities of bit size at most $b$.
@@ 2.1 Streaming algorithms
We only consider streaming problems where the input values are rational probabilities and there is a set of allowed rational output probabilties (this also covers the case, where the output is either 00 or $1$). In our context, the output probabilities will be approximations of some rational probability that is uniquely determined by the input sequence.
@@
A deterministic streaming algorithm can be formalized by a family of deterministic finite automata $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$, where the automaton $\mathcal{A}_{n,b}$ is responsible for inputs from the slice $\mathbb{P}_{b}^{\leq n}$. Hence, the input alphabet of $\mathcal{A}_{n,b}$ is $\mathbb{P}_{b}$. Let $Q_{n,b}$ be the set of states of $\mathcal{A}_{n,b}$. Its elements can be seen as the memory states of the streaming algorithm. There is a distinguished initial state in $\mathcal{A}_{n,b}$. The dynamics of the streaming algorithm is described by the state transition function $\delta_{n,b}:Q_{n,b}\times\mathbb{P}_{b}\to Q_{n,b}$ such that for a memory state $s\in Q_{n,b}$ and a new data value $q\in\mathbb{P}_{b}$, $\delta_{n,b}(s,q)\in Q_{n,b}$ is the new memory state. In addition, $\mathcal{A}_{n,b}$ has an output function $o_{n,b}:Q_{n,b}\to\mathbb{P}$ instead of a set of final states. Such an automaton computes a function $f_{\mathcal{A}_{n,b}}:\mathbb{P}_{b}^{\leq n}\to\mathbb{P}$ as follows: Starting from the initial state, the automaton reads the input stream $w\in\mathbb{P}_{b}^{*}$. After reading the last input number from $w$, the automaton is in a certain state $s\in Q_{n,b}$. The output value $f_{\mathcal{A}_{n,b}}(w)$ is then $o_{n,b}(s)$. A deterministic streaming algorithm $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy $f_{\mathcal{A}_{n,b}}(w)\in F(w)$ for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$. Finally, the space complexity of the streaming algorithm $\mathcal{D}$ is the mapping $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$, which is the number of bits required to encode the states in binary notation.
@@
Some comments regarding our definition are appropriate. First, our definition of a deterministic streaming algorithm is non-uniform in the sense that for every input slice $\mathbb{P}_{b}^{\leq n}$ we have a separate algorithm (the automaton $\mathcal{A}_{n,b}$). In principle the function $(n,b)\mapsto\mathcal{A}_{n,b}$ may be not computable. In a real streaming algorithm, this function will be certainly computable. On the other hand, the main focus in this paper is on lower bounds, and lower bounds for non-uniform algorithms also hold for uniform algorithms. Related to this point is the fact that we ignore in our definition of the space complexity $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$ the space needed for computing the the mapping $\delta_{n,b}$, i.e., the space needed to compute the next memory state from the previous memory state and the input data value (this space is what we called the space for internal computations in footnote 2). In general, this space will also grow with $n$ and $b$. The justification for ignoring this space is the same as for non-uniformity. A lower bound for our measure of space complexity also holds for the more realistic measure that incorporates the internal space needed for computing the state transformation mapping. Another point is that lower bound techniques for streaming algorithms (which are typically based on information theoretic arguments or communication complexity) only yield lower bounds on our space complexity.
@@
Randomized streaming algorithms can be defined analogously to deterministic streaming algorithms by taking for every $\mathcal{A}_{n,b}$ a probabilistic finite automaton (see for details on probabilistic automata). For an input stream $w\in\mathbb{P}_{b}^{*}$, we then obtain a random variable $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}$, whose support is the set of all output values $o_{n,b}(s)$ for $s\in Q_{s,b}$. For an output value $q$, $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}(q)$ is the probability that on input $w$ the probabilistic automaton $\mathcal{A}_{n,b}$ returns $q$. A randomized streaming algorithm $\mathcal{R}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy
$\mathsf{Prob}[\mathsf{R}_{\mathcal{A}_{n,b}}^{w}\in F(w)]\geq\frac{2}{3}$
for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$. The probability $\frac{2}{3}$ can be replaced by any $1-\lambda$ by running $\mathcal{O}(\log(1/\lambda))$ many independent copies of $\mathcal{A}_{n,b}$.
@@
We consider the following streaming problems in this paper, where $10\leq\epsilon<$ 1 is an approximation ratio and $q_{1},q_{2},\ldots,q_{n}\in\mathbb{P}$:
  • The streaming problem $\mathsf{APP}_{\epsilon}$ (approximation of products of probabilities) is
@@ 2.2 Communication complexity
Lower bounds for randomized streaming algorithms are typically proved by using lower bounds on randomized communication complexity; see [4, 8] for a detailed introduction. We only use the one-way setting where Alice sends a single message to Bob.
Consider a function $f\colon X\times Y\to\{0,1\}$ for some finite sets $X$ and $Y$, which we call a communication problem. A randomized one-way (communication) protocol $P$ for $f$ consists of two parties called Alice and Bob. The input for Alice (resp., Bob) is an element $x\in X$ and a random choice $r\in R$ (resp., $y\in Y$ and a random choice $s\in S$). Here, $R$ and $S$ are finite sets and the random choices of Alice and Bob are independent. The goal of Alice and Bob is to compute $f(x,y)$. For this, Alice computes from here input $x\in X$ and random choice $r\in R$ a message $a(x,r)\in\{0,1\}^{*}$ which is send to Bob. Bob then computes from $a(x,r)$, his input $y\in Y$ and random choice $s\in S$ the final output $b(a(x,r),y,s)\in\{0,1\}$ of the protocol. We also assume a probability distribution on the set $R$ (resp., $S$) of Alice’s (resp., Bob’s) random choices. The protocol $P$ computes $f$ if for all $(x,y)\in X\times Y$ we have
@@
The cost of the protocol is the maximum length $|a(x,r)|$ taken over all $(x,r)\in X\times R$. The randomized one-way communication complexity of $f$ is the minimal cost among all (one-way) randomized protocols that compute $f$. Here, the size of the finite sets $R$ and $S$ is not restricted. The choice of the constant $2/3$ in (1) is arbitrary in the sense that changing the constant to any $1-\lambda$ only changes the communication complexity by a constant (depending on $\lambda$), see [4, p. 30]. Also note that we only use the private version of randomized communication protocols, where Alice and Bob make private random choices from the sets $R$ and $S$, respectively, and their choices are not known to the other party (in contrast to the public version of randomized communication protocols).
@@
  • Bob’s input is a number $b\in\{1,\dots,m\}$.
  • Bob’s goal is to determine whether $a>b$.
The index-greater-than problem $\mathsf{IndexGT}_{m,n}$ is defined as follows:
@@
  • Bob’s input is a number $b\in\{1,\dots,m\}$ and a number $i\in\{1,\ldots,n\}$.
  • Bob’s goal is to determine whether $a_{i}<b$.
Theorem 2.1.
@@
  • The randomized one-way communication complexity of $\mathsf{GT}_{m}$ is $\Theta(\log m)$ [5].
  • The randomized one-way communication complexity of $\mathsf{IndexGT}_{m,n}$ is $\Theta(n\cdot\log m)$ [3].

### v2

@@ Abstract
We consider streaming algorithms for approximating a product of input probabilities up to multiplicative error of $1-\epsilon$. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(\log n+\log b-\log\epsilon-\Theta(1))$, where $n$ is length of the input stream and $b$ is the bit length of the input numbers. This matches an upper bound from Alur et al. up to a constant multiplicative factor. Moreover, we consider the threshold problem, where it is asked whether the product of the input probabilities is below a given threshold. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(n\cdot b)$. Finally, also the sliding window variant of the approximation problem for a product of input probabilities is considered.
@@ 1 Introduction
Our first result is a lower bound that matches the upper bound $2\log n+\log b-\log\epsilon$ up to a constant factor. This is also true for randomized streaming algorithms that produce a $(1-\epsilon)$-approximation of the product $\prod_{1\leq i\leq n}q_{i}$ with high probability. Our lower bound only holds if $\log(1/\epsilon)\leq b<\log(1/(1-\epsilon))\cdot n/2$. This is a reasonable setting. If we consider the approximation parameter $\epsilon$ as a small constant then the requirement becomes $\Omega(1)\leq b\leq\mathcal{O}(n)$. Since the length $n$ of the stream is usually very large, it is reasonable to assume $b\leq\mathcal{O}(n)$. In a realistic scenario one would even require $b\leq\mathcal{O}(\log n)$.
Our second result is a new lower bound for the threshold problem: the input consists again of a stream of probabilities $q_{1},q_{2},\ldots,q_{n}$, where $q_{i}=r_{i}/s_{i}$ and the integers $r_{i}$ and $s_{i}$ need at most $b$ bits. In addition, there is a threshold probability $t=r/s$, where $r$ and $s$ also need at most $b$ bits. The question is whether $\prod_{1\leq i\leq n}q_{i}<t$ holds. It was shown in [2, Theorem A.1] that every deterministic streaming algorithm for the general threshold problem, where the $q_{i}$ and $t$ are not restricted to the interval $[0,1]$, has to store $\Omega(n)$ many bits when $b=\log n$. It was asked whether this lower bound also holds when all $q_{i}$ and $t$ are probabilities. We prove that this is indeed the case and moreover improve the lower bound from $\Omega(n)$ to $\Omega(n\cdot b)$ if $b\geq\Omega(\log n)$ for randomized streaming algorithms. Note that the threshold problem can be always solved in space $\Omega(n\cdot b)$ by naively computing the product $\prod_{1\leq i\leq n}q_{i}$. On the other hand, if $n$ is very large compared to $b$, then there is a better (deterministic) algorithm that only needs $\mathcal{O}(2^{b})$ bits; see Theorem 4.1.
In the final section we consider the product approximation problem for rational probabilities in the sliding window setting. Here, the goal is to maintain a $(1-\epsilon)$-approximation of the product of the last $m$ probabilities from the input stream. The parameter $m$ is called the window size. We provide a space lower bound of $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$ for randomized sliding window algorithms, where as before $b$ is the bit length of the input probabilities. Currently, we are not able to match this lower bound by an upper bound. In Section 5 we provide two incomparable upper bounds: $2mb$ (which is trivial) and $m\cdot(\log m+\log b-\log\epsilon)$.
1.1 Related work
The first paper that studied the sliding window model formally was [5]. It is shown there that for a stream of positive integers of bit length $b$, a $(1-\epsilon)$-approximation of the sum of the numbers in a sliding window of length $m$ can be maintained in space $\mathcal{O}((1/\epsilon)(\log m+b)\log m)$. A matching lower bound is shown as well. A survey on sliding window algorithm can be found in [1].
@@ 2 Preliminaries
For a possibly infinite set $A$ we denote with $A^{*}$ the set of all finite words (or sequences) over the set $A$. We write $A^{\leq n}$ for the subset of words of length at most $n$. With $\log$ we always denote the logarithm to the base two. For a positive rational number $q=r/s$ with $r,s\in\mathbb{N}$ and $\mathsf{gcd}(r,s)=1$ we define the bit size of $q$ as $\|q\|=\max\{\lceil\log r\rceil,\lceil\log s\rceil\}$. In addition, we define $\|0\|=0$. Let $\mathbb{P}=\{q\in\mathbb{Q}:0\leq q\leq 1\}$ be the set of rational probabilities and $\mathbb{P}_{b}=\{q\in\mathbb{P}:\|q\|\leq b\}$ be the set of rational probabilities of bit size at most $b$.
@@ 2.1 Streaming algorithms
We only consider streaming problems where the input values are rational probabilities and there is a set of allowed rational output probabilties (this also covers the case, where the output is either 0 or $1$). In our context, the set of allowed output probabilities consists of all approximations of some rational probability that is uniquely determined by the input stream.
@@
A deterministic streaming algorithm can be formalized by a family of deterministic finite automata $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$, where the automaton $\mathcal{A}_{n,b}$ is responsible for inputs from the slice $\mathbb{P}_{b}^{\leq n}$. Hence, the input alphabet of $\mathcal{A}_{n,b}$ is $\mathbb{P}_{b}$. Let $Q_{n,b}$ be the set of states of $\mathcal{A}_{n,b}$. Its elements can be seen as the memory states of the streaming algorithm. There is a distinguished initial memory state $s_{0,n,b}\in Q_{n,b}$. The dynamics of the streaming algorithm is described by the state transition function $\delta_{n,b}:Q_{n,b}\times\mathbb{P}_{b}\to Q_{n,b}$ such that for a memory state $s\in Q_{n,b}$ and an incoming data value $q\in\mathbb{P}_{b}$, $\delta_{n,b}(s,q)\in Q_{n,b}$ is the new memory state. In addition, $\mathcal{A}_{n,b}$ has an output function $o_{n,b}:Q_{n,b}\to\mathbb{P}$ instead of a set of final states. The automaton $\mathcal{A}_{n,b}$ computes a function $f_{\mathcal{A}_{n,b}}:\mathbb{P}_{b}^{\leq n}\to\mathbb{P}$ as follows: Starting from the initial state, the automaton reads the input stream $w\in\mathbb{P}_{b}^{*}$. After reading the last input number from $w$, the automaton is in a certain state $s\in Q_{n,b}$. The output value $f_{\mathcal{A}_{n,b}}(w)$ is then $o_{n,b}(s)$. A deterministic streaming algorithm $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy $f_{\mathcal{A}_{n,b}}(w)\in F(w)$ for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$. Finally, the space complexity of the streaming algorithm $\mathcal{D}$ is the mapping $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$, which is the number of bits required to encode the states in binary notation. It is therefore the number of bits that $\mathcal{A}_{n,b}$ has to store.
@@
Some comments regarding our definition are appropriate. First, our definition of a deterministic streaming algorithm is non-uniform in the sense that for every input slice $\mathbb{P}_{b}^{\leq n}$ we have a separate algorithm (the automaton $\mathcal{A}_{n,b}$). In principle the function $(n,b)\mapsto\mathcal{A}_{n,b}$ may be not computable. In a real streaming algorithm, this function will be certainly computable. On the other hand, the main focus in this paper is on lower bounds, and lower bounds for non-uniform algorithms also hold for uniform algorithms. The few streaming algorithms that we present in this paper are moreover clearly uniform.
Related to the previous point is the fact that we ignore in our definition of the space complexity $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$ the space needed for computing the mapping $\delta_{n,b}$, i.e., the space needed to compute the next memory state from the previous memory state and the input data value. This space is what we called the space for internal computations in footnote 2. In general, this internal space will also grow with $n$ and $b$. In a non-uniform streaming algorithm, the function $(n,b,q,s)\mapsto\delta_{n,b}(q,s)$ might be even uncomputable. The justification for ignoring the space for internal computation is the same as for non-uniformity. A lower bound for our measure of space complexity also holds for the more realistic measure that incorporates the internal space needed for computing the state transformation mapping. Another point is that lower bound techniques for streaming algorithms (which are typically based on information theoretic arguments or communication complexity) can only yield lower bounds on our measure of space complexity: the classical application of communication lower bounds to streaming lower bounds constructs from a streaming algorithm (for a certain streaming problem) a communication protocol (for a certain communication problem). Thereby the messages exchanged between Alice to Bob are memory states of the streaming algorithm. Hence, a lower bound for the communication problem yields a lower bound for the bit length of the memory states of the streaming algorithm. This line of argument does not take the memory for internal computations in the streaming algorithm into account.
@@
Randomized streaming algorithms can be defined analogously to deterministic streaming algorithms by taking for every $\mathcal{A}_{n,b}$ a probabilistic finite automaton (see [9] for details on probabilistic automata). For an input stream $w\in\mathbb{P}_{b}^{*}$, we then obtain a random variable $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}$, whose support is the set of all output values $o_{n,b}(s)$ for $s\in Q_{n,b}$. For an output value $q\in\mathbb{P}$, $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}(q)$ is the probability that on input $w$ the probabilistic automaton $\mathcal{A}_{n,b}$ returns $q$. A randomized streaming algorithm $\mathcal{R}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy
$\mathsf{Prob}[\mathsf{R}_{\mathcal{A}_{n,b}}^{w}\in F(w)]\geq\frac{2}{3}$ (1)
for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$.
@@
We consider the following streaming problems in this paper, where $0\leq\epsilon<1$ is an approximation ratio and $q_{1},q_{2},\ldots,q_{n}\in\mathbb{P}$:
  • The approximation of products of probabilities, $\mathsf{APP}_{\epsilon}$ for short, is
@@
It is easy to see that for these streaming problems the error probability $\leq\frac{1}{3}$ from (1) can be reduced to any $\lambda<1/3$ by running $\mathcal{O}(\log(1/\lambda))$ many independent copies of $\mathcal{A}_{n,b}$. In the case of $\mathsf{APP}_{\epsilon}$, the output value will be the median of the output values of the independent copies of $\mathcal{A}_{n,b}$, whereas for $\mathsf{TPP}$ (where the output is boolean) the output bit is obtained by a majority vote.
@@ 2.2 Communication complexity
Lower bounds for randomized streaming algorithms are typically derived from lower bounds on randomized communication complexity [6, 11]. We only use the one-way setting, where Alice sends a single message to Bob.
Consider a function $f\colon X\times Y\to\{0,1\}$ for some finite sets $X$ and $Y$, which we call a communication problem. A randomized one-way (communication) protocol $P$ for $f$ consists of two parties called Alice and Bob. The input for Alice (resp., Bob) is an element $x\in X$ and a random choice $r\in R$ (resp., $y\in Y$ and a random choice $s\in S$). Here, $R$ and $S$ are finite sets and the random choices of Alice and Bob are independent. The goal of Alice and Bob is to compute $f(x,y)$ with high probability. For this, Alice computes from here input $x\in X$ and random choice $r\in R$ a message $a(x,r)\in\{0,1\}^{*}$ and sends it to Bob. Bob then computes from $a(x,r)$, his input $y\in Y$ and random choice $s\in S$ the final output $b(a(x,r),y,s)\in\{0,1\}$ of the protocol. We also assume a probability distribution on the set $R$ (resp., $S$) of Alice’s (resp., Bob’s) random choices. The protocol $P$ computes $f$ if for all $(x,y)\in X\times Y$ we have
@@
The cost of the protocol is the maximum length of $a(x,r)$ taken over all $(x,r)\in X\times R$. The randomized one-way communication complexity of $f$ is the minimal cost among all (one-way) randomized protocols that compute $f$. Here, the size of the finite sets $R$ and $S$ is not restricted. The choice of the constant $2/3$ in (2) is arbitrary in the sense that changing the constant to any $1-\lambda$ only changes the communication complexity by a constant (depending on $\lambda$), see [6, p. 30]. Also note that we only use the private version of randomized communication protocols, where Alice and Bob make private random choices from the sets $R$ and $S$, respectively, and their choices are not known to the other party (in contrast to the public version of randomized communication protocols).
@@
  • Bob’s input is a number $a^{\prime}\in\{1,\dots,m\}$.
  • Bob’s goal is to determine whether $a>a^{\prime}$.
The index-greater-than problem $\mathsf{IGT}_{m,n}$ is defined as follows:
@@
  • Bob’s input is a number $a^{\prime}\in\{1,\dots,m\}$ and a number $i\in\{1,\ldots,n\}$.
  • Bob’s goal is to determine whether $a_{i}>a^{\prime}$.
Theorem 2.1
@@
  • $\mathsf{GT}_{m}$ has randomized one-way communication complexity $\Theta(\log m)$ [7].
  • $\mathsf{IGT}_{m,n}$ has randomized one-way communication complexity $\Theta(n\cdot\log m)$ [4].